ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 227, 2018

Adams, F. and J. Zapletal Cardinal invariants of closed graphs  861

Balletti, G. and A. Higashitani Universal inequalities in Ehrhart Theory

843

Barnea, I. and S. Shelah The abelianization of inverse limits of groups

455

Beeker, B. and N. Lazarovich Stallings’ folds for cube complexes  331

Beltrán, A., M. J. Felipe and C. Melchor Multiplying a conjugacy class by its inverse in a finite group

811

Berglund, A. and K. Hess Homotopical Morita theory for corings  239
Bodin, A. and M. Borodzik Intermediate links of plane curves  63

Borodzik, M. See A. Bodin Bravo, M. and R. Cominetti Sharp convergence rates for averaged

nonexpansive maps

163

Brian, W. R. and P. Oprocha Ultrafilters and Ramsey-type shadowing phenomena in topological dynamics

423

Butler, C. Measurable rigidity of the cohomological equation for linear cocycles over hyperbolic systems

27

Bywaters, T. P., H. Glöckner and S. Tornier Contraction groups and passage to subgroups and quotients for endomorphisms of totally disconnected locally compact groups

691

Carnevale, A., S. Shechter and C. Voll Enumerating traceless matrices over compact discrete valuation rings

957

Chebotar, M. On differential polynomial rings over locally nilpotent rings

233

Clark, P. L. and P. Pollack Pursuing polynomial bounds on torsion

889

Cominetti, R. See M. Bravo Cristea, M. On the lightness of the mappings satisfying generalized

inverse modular inequalities

545

Di Plinio, F. and I. Parissis A sharp estimate for the Hilbert transform along finite order lacunary sets of directions

189

Feldheim, N. D. Variance of the number of zeroes of shift-invariant Gaussian analytic functions

753

Felipe, M. J. See A. Beltrán Figueiredo, G. M., J. R. Santos Júnior and A. Suárez

Structure of the set of positive solutions of a non-linear Schrödinger equation

485

Glöckner, H. See T. P. Bywaters Goldbring, I., B. Hart and H. Towsner Explicit sentences

distinguishing McDuff’s II_{1} factors

365

Hart, B. See I. Goldbring Haviv, I. and D. Levy Symmetric complete sum-free sets in cyclic

groups

931

Heittokangas, J., K. Ishizaki, K. Tohge and Z.-T. Wen Zero distribution and division results for exponential polynomials

397

Hess, K. See A. Berglund Higashitani, A. See G. Balletti Huang, B., S. Liu and Z. Xu Mollification and non-vanishing of

automorphic L-functions on GL(3)

597

Ishizaki, K. See J. Heittokangas Jambor, S., A. Litterick and C. Marion On finite simple images

of triangle groups

131

Lazarovich, N. See B. Beeker Levy, D. See I. Haviv Li, G. and X. Wang Fibred cofinitely-coarse embeddability of box

families and proper isometric affine actions on uniformly convex Banach spaces

793

Litterick, A. See S. Jambor Liu, S. See B. Huang Lu, F. See C. Yang Lüdtke, M. A birational anabelian reconstruction theorem for curves

over algebraically closed fields in Arbitrary Characteristic

987

Marion, C. See S. Jambor Marquis, L. and J. Souto Surface groups of diffeomorphisms of the

interval

379

Melchor, C. See A. Beltrán

Miller, J. S. and J. Rute Energy randomness  1
Odrzygóźdź, T. Cubulating random groups in the square model  623

Oprocha, P. See W. R. Brian Parissis, I. See F. Di Plinio Pollack, P. See P. L. Clark Rapaport, A. A dimension gap for continued fractions with indepen-

dent digits—the non stationary case 911 Raz, O. E., M. Sharir and F. de Zeeuw The Elekes–Szabó

Theorem in four dimensions

663

Rebrova, E. and K. Tikhomirov Coverings of random ellipsoids, and invertibility of matrices with i.i.d. heavy-tailed entries

507

Ribón, J. Finite dimensional groups of local diffeomorphisms  289

Rute, J. See J. S. Miller Santos Júnior, J. R. See G. M. Figueiredo Sharir, M. See O. E. Raz Shechter, S. See A. Carnevale Shelah, S. See I. Barnea Simon, R. S. and G. Tomkowicz A Bayesian Game without

ϵ-equilibria

215

Sklinos, R. The free group does not have the finite cover property  563

Souto, J. See L. Marquis Suárez, A. See G. M. Figueiredo Thom, A. and J. Wilson Some geometric properties of metric

ultraproducts of finite simple groups

113

Tikhomirov, K. See E. Rebrova Tohge, K. See J. Heittokangas Tomkowicz, G. See R. S. Simon Tornier, S. See T. P. Bywaters Towsner, H. See I. Goldbring Voll, C. See A. Carnevale Wang, X. See G. Li Wen, Z.-T. See J. Heittokangas Wilson, J. See A. Thom Xu, Z. See B. Huang

Yang, C. and F. Lu Lie isomorphisms of reflexive algebras. II  827

Zapletal, J. See F. Adams de Zeeuw, F. See O. E. Raz